two wires of a and b with circular cross section made up of the same material wih equal lengths.suppose Ra=3Rb,then what is the ratio of the radius of the 
wire a to that of b?

Grade:12

1 Answers

Arun
4 years ago
 
Dear student 
 
R1 = 3 R2
 
rho * l / A1 = 3 rho * l / A2
 
A2/ A1 = 3
 
(r1)^2 / (r2)^2 = 1/3
 
hence r1 / r2 = 1 / sqrt(3)